OpenStudy (anonymous):

“Think of a number. Multiply it by 20, divide the answer by 9, and then subtract seven from the result.” a. Make a table of values and plot the function that represents this sentence. b. If you picked 0 as your starting number, what number would you end up with? c. To end up with 12, what number would you have to start out with?

OpenStudy (the_fizicx99):

\(\ f(x) = \dfrac{x \times 20}{9} - 7 \) Make a table by plugging in -2,-1,0,1,2,3 and so on. This will help you plot it on a graph. Then replace f(x) for y and plot it into a graph. When you plot it at 0 you'll find the starting point, often called the y intercept or principle. For the last part you can plot it and see which x value will give you 12 :> OR 12 = \(\ \dfrac{x \times 20}{9} - 7 add both sides by 7 9(19) = 9(x * 20)/9 171 = x * 20 /20 8.55 = x so when x is 8.55 it'll give you 12

OpenStudy (the_fizicx99):

Well that was a fail X___x 12 = \(\ \dfrac{x \times 20}{9} - 7 \)

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

thanks for your help

OpenStudy (the_fizicx99):

Np, gl